# System Identification for Dynamic Modeling of a Bumper Car

### v1

@@ 1 Introduction
Autonomous driving has emerged as a transformative technology, attracting significant attention due to its potential to enhance road safety. Despite consistent progress and strong evidence supporting its safety, public acceptance of autonomous driving remains limited. For instance, the comparative study by Lillo et al. (2023) demonstrated that Waymo’s autonomous service achieved markedly fewer injury and property damage claims per mile than human drivers, highlighting its superior safety performance. Nevertheless, societal concerns and limited exposure continue to hinder trust in this technology. Motivated by these observations, we aim at building an experimental platform where autonomous driving safety can be intuitively and convincingly demonstrated. We develop an autonomous bumper car platform that showcases collision avoidance in a playful and interactive manner. Within this setup, visitors can directly engage with the system, attempting to provoke collisions while the vehicle autonomously reacts to prevent them.
@@
A natural starting point is the family of planar bicycle models, which balance modeling fidelity and computational cost for path-planning and prediction tasks (Jazar, 2008). Two variants are very popular in the literature: the dynamic bicycle model, which is derived from the Newton–Euler formalism and explicitly accounts for tire-force interactions, and the kinematic bicycle model, which is based on the geometric configuration of the vehicle chassis.
@@
1.1 Related Work
By considering vehicles dynamics modeling, recent work extends beyond purely first-principles modeling by incorporating data-driven identification within white/grey/black-box pipelines. A typical field of application of such dedicated data-driven models is autonomous racing, where the main control challenge arises from pushing the vehicle to the handling limits. In the following paragraphs, we review relevant works that embody and advance these methodologies.
While Newton-Euler models are close to physical principles, the underlying tire model brings its highest uncertainty. Thus, the parameters of the Magic Formula in Pacejka (2005) are estimated online using a Long Short-Term Memory (LSTM) model in Kim et al. (2022) and Chrosniak et al. (2024). In Paparazzo et al. (2025) the Newton-Euler parameters are updated online using constrained Sparse Identification of Nonlinear Dynamics (SINDy) on recent data, allowing the autonomous system to improve trajectory tracking in an MPC-framework. A linearization of the Newton-Euler model is combined with a Recurrent Neural Network (RNN) residual model in Chen et al. (2024), compensating for nonlinear lateral dynamics.
@@ 1.1 Related Work
Previous works rely on assumptions that are not always met by high-maneuverable vehicles. First, their reported accuracy can degrade in regimes with advanced steering range and pronounced slip, as encountered on our platform. Second, while Newton–Euler formulations can represent these regimes, their numerical stiffness and integration cost can be prohibitive in real-time applications. To the best of our knowledge, none of the existing models retains the accuracy of high steering Newton–Euler dynamics while achieving kinematic-like computational speed suitable for embedded deployment.
1.2 Contributions
@@
2) We adapt the extended kinematic formulation of Ghosn et al. (2024) to bumper cars with extended steering, closely tying the system state to the system actuation. On top of the kinematic backbone, we employ the Sparse Identification of Nonlinear Dynamical Systems (SINDy) framework to capture the vehicle state evolution in a physics-consistent manner and Multilayer Perceptron (MLP) neural networks to model the remaining dynamics. Furthermore, through a progressive transition from black-box to grey-box representations, we examine the advantages of embedding physical structure into the learned model.
@@ 2 Bumper Car Models
The bumper car consists a vehicle body with one front wheel and two rear wheels. The front wheel is both steered and motorized, while the rear axle comprises two non-driven wheels. Each component of the bumper car, as depicted in Figure 1, is briefly discussed.
@@
with maximum steer rate $d>0$ and time constant $T_{\mathrm{s}}>0$.
@@
In the following, we introduce both models and describe the modifications made to address the previously discussed limitations. The derivation and notation employed for both models closely follow the conventions established in Jazar (2008), specifically regarding tire slip angle $\alpha_{i}$, tire heading angle $\beta_{i}$ and wheelbase distance $l_{i}$ with subscript $i$ denoting the front (f) or rear (r) axle (cf. Figure 3 and 4).
2.1 Newton-Euler model
@@
$\dot{v}_{\mathrm{x}}$ $=F_{\mathrm{x}}/{m}+\omega\,v_{y}$ (2)
$\dot{v}_{\mathrm{y}}$ $={F_{\mathrm{y}}}/{m}-\omega\,v_{\mathrm{x}}$
$\dot{\omega}$ $={M_{\mathrm{z}}}/{I_{\mathrm{z}}}$
where $m$ is the mass, $I_{z}$ the moment of inertia and $F_{\mathrm{x}}$, $F_{\mathrm{y}}$ and $M_{\mathrm{z}}$ the expressions for the longitudinal force, lateral force, and yaw moment.
The model accuracy and numerical well-posedness primarily depend on two design aspects: the selection and mathematical formulation of the tire forces, and the implementation of the tire slip angle computation.
We first address the choice of tire forces and derive
@@
based on the lateral slip forces $F_{\alpha,i}$ and the longitudinal drag and motor forces $F_{\mathrm{d}}$ and $F_{\mathrm{m}}$, as illustrated in Figure 3.
@@
While longitudinal tire force is neglected at the rear wheel, we model drag-induced force at the front tire as a second order polynomial
$F_{d}=R_{1}\,v_{\mathrm{f,x}}+R_{2}\,v_{\mathrm{f,x}}^{2}$ (4)
with drag coefficients $R_{1}$ and $R_{2}$.
For a bumper car, which generally operates at low speeds and small slip angles, a linear tire force model approximates most of the dynamics, as it captures the proportional relation $F_{\mathrm{\alpha},i}\propto\alpha_{i}$ in that operating range. Using models with many parameters—such as the Magic Formula—is not practical because those parameters cannot be reliably identified from datasets with limited high-slip tire behavior.
To balance simplicity and realism, a suitable compromise (Zhang and Liu, 2018) is the following piecewise linear tire model
@@
where $C_{\alpha,i}$ is the cornering stiffness and $\hat{F}_{\alpha,i}$ the maximum lateral tire force of axis $i\in\{f,r\}$, respectively.
The second design aspect of the Newton-Euler model concerns the numerical robustness of tire slip computation, when the velocity vector approaches zero (i.e., the vehicle approaches standstill) and the slip angle $\alpha_{i}$ loses physical meaning. Kim et al. (2018) introduced an Advanced Slip Ratio (ASR), where a lower bound is imposed on the longitudinal velocity when computing the tire slip angle, thus implying $\alpha\to 0$ for low velocities.
Moreover, the extended steering range induces negative longitudinal velocity, causing the rear wheel to rotate in reverse. As a result, the rear tire slip is given by $\alpha_{\mathrm{r}}=\beta_{\mathrm{r}}-\delta_{\mathrm{r}}$ where $\delta_{\mathrm{r}}$ is 0° if the wheel spins forward and 180° if backward. Due to the trigonometric properties of the $\arctan(\cdot)$ function, we compute the tire heading angle using the ASR and the absolute value of the longitudinal tire velocity as
Kim et al. (2018) shows that the integration of ASR limits the lateral tire dynamics, however, the lateral tire dynamics remain faster than the dynamics of the vehicle body inertia, making the Newton-Euler model stiff. Consequently, discretization ($T=$0.1\text{\,}\mathrm{s}$\$) performed using RK4 with two subsequent steps yields the body frame velocity evolution as
@@ 2.2 Kinematic model
Secondly, the assumption of slip-free movement does not hold in our context. Therefore, we derive the kinematic formulation considering the tire slip, similar to Ghosn et al. (2024). The kinematic bicycle model is depicted in Figure 4
@@
with the corresponding set of equations
$v_{\mathrm{x}}$ $=\cos{\beta_{\mathrm{f}}}\,v_{\mathrm{f}}$ (9)
$\text{where}\;\;\beta_{\mathrm{f}}=\delta+\alpha_{\mathrm{f}},\quad\beta_{\mathrm{r}}=\alpha_{\mathrm{r}}\,.$
which describe the static output function
$\bm{v}_{k}=g_{\mathrm{kin}}(\bm{x}_{k})\quad\text{with}\quad\bm{x}_{k}=\begin{bmatrix}v_{\mathrm{f}}&\alpha_{\mathrm{f}}&\alpha_{\mathrm{r}}&\delta\end{bmatrix}^{T}\,.$ (10)
The evolution of $\delta$ follows from the time discretization of (1). The dynamics of the remaining state $\begin{bmatrix}v_{\mathrm{f}}&\alpha_{\mathrm{f}}&\alpha_{\mathrm{r}}\end{bmatrix}^{T}$ are not explicitly characterized from first principles. Rather, they are assumed to follow the general discrete-time model representation:
@@ 3 Data-Driven modeling
Secondly, black-box approaches using multilayer perceptrons (MLPs) are investigated. We will show that these models are capable of capturing complex nonlinearities but suffer from limited physical interpretability and sensitivity to data quality, which complicates reliable online adaptation. Technical details for the implementation are reported in Appendix C.
Eventually, grey-box methods are employed to merge the advantages of both approaches. By incorporating physical knowledge within a kinematic bicycle framework through Sparse Identification of Nonlinear Dynamical Systems (SINDy), an explicit and interpretable model structure is obtained. Finally, we propose a hybrid residual modeling approach that combines the strengths of physics-based models and neural networks, enhancing prediction accuracy while preserving interpretability.
@@
3.1 White-box modeling
The geometric parameters of the Newton–Euler model ($l_{\mathrm{f}}$, $l_{\mathrm{r}}$, $m$, $\hat{\delta}$) and parameters related to the actuators ($d$, $T_{\mathrm{s}}$) can be directly measured or obtained from hardware specifications. However, a subset of model parameters are unknown or difficult to measure accurately. Therefore, they are estimated with a genetic algorithm due to the highly nonlinear and partially non-differentiable dependence of the model outputs on its parameters (Goldberg, 1991). The training is performed by minimizing the NMSE on the one-step body frame velocity predictions, namely setting $\hat{\bm{q}}=f_{\mathrm{NE}}(\bm{v},\bm{u})$ and $\bm{q}=\bm{v}$ in (12). The resulting parameters are reported in Appendix B.
3.2 Black-box modeling
@@ 3.3 Grey-box modeling
The function library $\bm{\Theta}^{T}(\bm{x}_{k},\bm{u}_{k})$ is constructed based on available physical insight: generic polynomial terms are used when prior knowledge is limited, whereas trigonometric or physically motivated functions can be included to capture expected behaviors such as periodicity or symmetry. The coefficient vector $\bm{\Xi}$ contains the weights assigned to each candidate function, and sparsity constraints are applied to identify only the few terms that significantly contribute to the system dynamics (de Silva et al., 2020; Kaptanoglu et al., 2022). The candidate libraries $\bm{\Theta}_{i}$ consisting of entries $q_{i,j}$ are derived as follows.
@@
leverages the physics-informed library described above to produce interpretable representations. The resulting state equation is reported in Appendix D.
@@
The training is performed on the NMSE of the one-step state prediction residual error $\bm{e}_{\mathrm{k+1}}=\tilde{\bm{x}}_{k+1}-f_{\mathrm{SINDy}}({\bm{x}}_{k},\bm{u}_{k})$, namely setting $\hat{\bm{q}}=f_{\mathrm{RES-MLP}}(\bm{x}_{k},\bm{u}_{k})$ and $\bm{q}=\bm{e}$ in (12).
@@ 4 Experimental validation
All data were recorded in a rectangular bumper-car arena of approximately $7\times 12\,$\mathrm{m}$\$. The bumper car pose $\bm{p}$ was measured by an OptiTrack motion-capture system using eight cameras placed around the perimeter. The body–fixed velocities $\bm{v}$ or kinematic state $\bm{x}$ according to (10) are not directly measured and are therefore estimated with an Extended Kalman Filter (EKF) described in Appendix A. We recorded slightly over one hour of driving at $10\text{\,}\mathrm{Hz}$, covering a wide range of operating scenarios. The dataset was then split into validation data ($110\text{\,}\mathrm{s}$) and training data (the remaining part).
The comparison metrics associated with the models presented in Section 3 are reported in Table 1. We evaluate each model by initializing it with the starting sample of the validation dataset and driving it with the validation input sequence. By comparing the resulting predicted trajectory against the actual validation trajectory, we assess performance in terms of the NMSE and the total computational time required for the simulation.
@@
To conclude this analysis, we compare the Newton–Euler and the K-SINDy-MLP models to highlight the improvements obtained through data-driven modeling. Figure 5 and Figure 6 illustrate a validation trajectory extracted from the dataset and the corresponding predicted trajectories obtained with the two models. At several points along the trajectories, marked by crosses, the predicted state is reinitialized to the true validation state to test different initial conditions. A close-up view between 66.4 and 68.4 seconds captures the vehicle’s transition from backward motion (steering angle exceeding 90°) to forward motion (steering angle below 90°). Within this critical segment, it is evident that the K-SINDy-MLP model substantially outperforms the Newton–Euler model (as well as its linearized version, see Figure 2) in accurately predicting the vehicle’s state, highlighting the superior capacity of the hybrid approach to manage complex dynamical regimes.
@@
Figure 5: Comparison between the experimental vehicle trajectory and the predicted trajectory obtained with the Newton-Euler model. The model is reinitialized every two seconds to test its effectiveness in different scenarios.
@@
Figure 6: Comparison between the experimental vehicle trajectory and the predicted trajectory obtained with the K-SINDy-MLP model. The model is reinitialized every two seconds to test its effectiveness in different scenarios.

### v2

# System Identification for Dynamic Modeling of Large Steering Angle Vehicles

@@ 1 Introduction
Autonomous driving has emerged as a transformative technology, attracting significant attention due to its potential to enhance road safety. Despite consistent progress and strong evidence supporting its safety, public acceptance of autonomous driving remains limited. For instance, the comparative study by Lillo et al. (2023) demonstrated that Waymo’s autonomous service achieved markedly fewer injury and property damage claims per mile than human drivers, highlighting its superior safety performance. Nevertheless, societal concerns and limited exposure continue to hinder trust in this technology. Motivated by these observations, we aim at building an experimental platform where autonomous driving safety can be intuitively and convincingly demonstrated in a bumper car ride, showcasing collision avoidance in a playful and interactive manner. Within this setup, visitors can directly engage with the system, attempting to provoke collisions while the vehicle autonomously reacts to prevent them.
@@
A natural starting point is the family of planar bicycle models, which balance modeling fidelity and computational cost for path-planning and prediction tasks (Jazar, 2008). Two variants are popular in the literature: the dynamic bicycle model, which is derived from the Newton–Euler formalism and explicitly accounts for tire-force interactions, and the kinematic bicycle model, which is based on the geometric configuration of the vehicle chassis.
@@
Related Work
By considering vehicles dynamics modeling, recent works extend beyond purely first-principles modeling by incorporating data-driven identification within white/grey/black-box pipelines. A typical field of application of such dedicated data-driven models is autonomous racing, where the main control challenge arises from pushing the vehicle to the handling limits. In the following paragraphs, we review relevant works that embody and advance these methodologies.
While Newton-Euler models are close to physical principles, the underlying tire model brings its highest uncertainty. Thus, the parameters of the Magic Formula in Pacejka (2005) are estimated online using a Long Short-Term Memory (LSTM) model in Kim et al. (2022) and Chrosniak et al. (2024). In Paparazzo et al. (2025) the Newton-Euler parameters are updated online using constrained Sparse Identification of Nonlinear Dynamics (SINDy, Brunton et al. (2016)) on recent data, allowing the autonomous system to improve trajectory tracking in an MPC-framework. A linearization of the Newton-Euler model is combined with a Recurrent Neural Network (RNN) residual model in Chen et al. (2024), compensating for nonlinear lateral dynamics.
@@
Previous works rely on assumptions that are not always met by high-maneuverable vehicles. First, their reported accuracy can degrade in regimes with wide steering range and pronounced slip, as encountered on our platform. Second, while Newton–Euler formulations can represent these regimes, their numerical stiffness and integration cost can be prohibitive in real-time applications. To the best of our knowledge, none of the existing models retains the accuracy of high steering Newton–Euler dynamics while achieving kinematic-like computational speed suitable for embedded deployment.
Contributions
@@
2) We adapt the extended kinematic formulation of Ghosn et al. (2024) to bumper cars with extended steering, closely tying the system state to the system actuation. On top of the kinematic backbone, we employ the SINDy framework to capture the vehicle state evolution in a physics-consistent manner and Multilayer Perceptron (MLP) neural networks to model the remaining dynamics. Furthermore, through a progressive transition from black-box to grey-box representations, we examine the advantages of embedding physical structure into the learned model.
@@ 2 Physics Based Models
The bumper car consists a vehicle body with one front wheel and two rear wheels. The front wheel is both steered and motorized, while the rear axle comprises two non-driven wheels. Each component of the bumper car is depicted in Figure 1.
@@
The steering and motor dynamics are described by first-order models with saturation and a proportional motor controller, detailed in Appendix A.
@@
In the following, we introduce both models and describe the modifications made to address the previously discussed limitations. The derivation and notation employed for both models follow the conventions established in Jazar (2008), specifically regarding tire slip angle $\alpha_{i}$, tire heading angle $\beta_{i}$ and wheelbase distance $l_{i}$ with subscript $i\in\{f,r\}$ denoting the front (f) or rear (r) axle (cf. Figure 3 and 4).
Newton-Euler model
@@
$\dot{v}_{\mathrm{x}}=F_{\mathrm{x}}/{m}+\omega\,v_{y},\,\,\dot{v}_{\mathrm{y}}={F_{\mathrm{y}}}/{m}-\omega\,v_{\mathrm{x}},\,\,\dot{\omega}={M_{\mathrm{z}}}/{I_{\mathrm{z}}}$ (1)
where $m$ is the mass, $I_{z}$ the moment of inertia and $F_{\mathrm{x}}$, $F_{\mathrm{y}}$ and $M_{\mathrm{z}}$ the expressions for the longitudinal force, lateral force, and yaw moment. We further derive
@@
based on the lateral slip, longitudinal drag and motor forces $F_{\alpha,i}$, $F_{\mathrm{d}}$ and $F_{\mathrm{m}}$, as illustrated in Figure 3.
@@
While longitudinal tire force is neglected at the rear wheel, we model drag-induced force at the front tire $F_{d}=R_{1}\,v_{\mathrm{f,x}}+R_{2}\,v_{\mathrm{f,x}}^{2}$ with drag coefficients $R_{1}$ and $R_{2}$.
For a bumper car, which generally operates at low speeds and small slip angles, a linear tire force model approximates most of the dynamics, as it captures the proportional relation $F_{\mathrm{\alpha},i}\propto\alpha_{i}$ in that operating range. Using models with many parameters—such as the Magic Formula—is not practical because those parameters cannot be reliably identified from datasets with limited high-slip tire behavior. To balance simplicity and realism, a suitable compromise (Zhang and Liu, 2018) is the following piecewise linear tire model
@@
where $C_{\alpha,i}$ is the cornering stiffness and $\hat{F}_{\alpha,i}$ the maximum lateral tire force. High steering angles ($|\delta|>90^{\circ}$) induce negative longitudinal velocity of the rear wheel. At zero-crossing the slip angle  $\alpha_{i}$ loses physical meaning. To handle these situations, we adapt the Advanced Slip Ratio of Kim et al. (2018); additional details are reported in Appendix B. Discretization ($T=$0.1\text{\,}\mathrm{s}$\$) is performed using RK4 with two subsequent steps yields the body frame velocity evolution as
@@
Secondly, the assumption of slip-free movement does not hold in our context. Therefore, we derive the kinematic formulation considering the tire slip, similar to Ghosn et al. (2024). The kinematic bicycle model is depicted in Figure 4 with the static output function $\bm{v}_{k}=g_{\mathrm{kin}}(\bm{x}_{k})$ given by
$v_{\mathrm{x}}$ $=\cos{\beta_{\mathrm{f}}}\,v_{\mathrm{f}}$ (5)
based on the system state $\bm{x}_{k}=\begin{bmatrix}v_{\mathrm{f}}&\alpha_{\mathrm{f}}&\alpha_{\mathrm{r}}&\delta\end{bmatrix}^{T}$ with heading angles $\beta_{\mathrm{f}}=\delta+\alpha_{\mathrm{f}}$ and $\beta_{\mathrm{r}}=\alpha_{\mathrm{r}}$.
@@
The evolution of $\delta$ follows from the time discretization of the first-order steering dynamics. The dynamics of the remaining state $\tilde{\bm{x}}=\begin{bmatrix}v_{\mathrm{f}}&\alpha_{\mathrm{f}}&\alpha_{\mathrm{r}}\end{bmatrix}^{T}$ are not explicitly characterized from first principles. Rather, they are assumed to follow the general discrete-time model representation:
@@ 3 Data-Driven modeling
Secondly, black-box approaches using MLPs are investigated. We show that these models are capable of capturing complex nonlinearities but suffer from limited physical interpretability and sensitivity to data quality, which complicates reliable online adaptation. Technical details for the implementation are reported in Appendix C.
Eventually, grey-box methods are employed to merge the advantages of both approaches. By incorporating physical knowledge within a kinematic bicycle framework through SINDy, an explicit and interpretable model structure is obtained. Finally, we propose a hybrid residual modeling approach that combines the strengths of physics-based models and neural networks, enhancing prediction accuracy while preserving interpretability.
@@
White-box modeling
The geometric parameters of the Newton–Euler model ($l_{\mathrm{f}}$, $l_{\mathrm{r}}$, $m$, $\hat{\delta}$) and parameters related to the actuators ($d$, $T_{\mathrm{s}}$) can be directly measured or obtained from hardware specifications. The remaining parameters are estimated with a genetic algorithm due to the highly nonlinear and partially non-differentiable dependence of the model outputs on its parameters (Goldberg, 1991). The training is performed by minimizing the NMSE on the one-step body frame velocity predictions, namely setting $\hat{\bm{q}}=f_{\mathrm{NE}}(\bm{v},\bm{u})$ and $\bm{q}=\bm{v}$ in (7). The resulting parameters are reported in Appendix D.
Black-box modeling
@@
The function library $\bm{\Theta}^{T}(\bm{x}_{k},\bm{u}_{k})$ is constructed based on available physical insight: generic polynomial terms are used when prior knowledge is limited, whereas trigonometric or physically motivated functions can be included to capture expected behaviors such as periodicity or symmetry. The coefficient vector $\bm{\Xi}$ contains the weights assigned to each candidate function, and sparsity constraints are applied to identify only the few terms that significantly contribute to the system dynamics (de Silva et al., 2020). The candidate libraries $\bm{\Theta}_{i}$ consisting of entries $q_{i,j}$ are derived as follows.
@@
leverages the physics-informed library described above to produce interpretable representations. The resulting state equation is reported in Appendix E.
@@
The training of the MLP is performed on the NMSE of the one-step state prediction residual error $\bm{e}_{\mathrm{k+1}}=\tilde{\bm{x}}_{k+1}-f_{\mathrm{SINDy}}({\bm{x}}_{k},\bm{u}_{k})$, namely setting $\hat{\bm{q}}=f_{\mathrm{RES-MLP}}(\bm{x}_{k},\bm{u}_{k})$ and $\bm{q}=\bm{e}$ in (7).
@@ 4 Experimental validation
Data was recorded in a rectangular arena of $7\times 12\,$\mathrm{m}$\$ using an OptiTrack motion-capture system for vehicle pose $\bm{p}$. The body–fixed velocities $\bm{v}$ and kinematic state $\bm{x}$ are not directly measured and are therefore estimated with an Extended Kalman Filter (EKF) described in Appendix F. We recorded one hour of driving at $10\text{\,}\mathrm{Hz}$ and an additional sequence of $110\text{\,}\mathrm{s}$ for validation.
The comparison metrics associated with the derived models are reported in Table 1. We evaluate each model by initializing it with the starting sample of the validation dataset and driving it with the validation input sequence. By comparing the predicted and actual velocity, we assess performance in terms of the NMSE and the total computational time required for the simulation.
@@
To conclude this analysis, we compare the Newton–Euler and the K-SINDy-MLP models to highlight the improvements obtained through data-driven modeling. Figure 5(a) and Figure 5(b) illustrate an excerpt from the validation trajectory and the corresponding predicted trajectories. At several points along the trajectories, marked by crosses, the vehicle pose is reinitialized to mitigate integration error. A close-up view between 66.4 and 68.4 seconds captures the vehicle’s transition from backward motion ($|\delta|\!>\!90^{\circ}$) to forward motion ($|\delta|\!<\!90^{\circ}$). Within this critical segment, it is evident that the K-SINDy-MLP model substantially outperforms the Newton–Euler model (as well as its linearized version, see Figure 2), highlighting the superior capacity of the hybrid approach to manage complex dynamical regimes.
@@
(a) Newton-Euler model
@@
(b) K-SINDy-MLP model.
Figure 5: Comparison between the experimental vehicle trajectory and the predicted trajectory. The models are reinitialized every two seconds to test their effectiveness in different scenarios.